# Ultrastrongly-coupled and Directionally-nonreciprocal Magnon-polaritons in Magnetochiral Metamolecules

### v2

@@ Abstract
We report an experimental demonstration of magnon-polaritons in a three-dimensional metamolecule lacking time-reversal and space-inversion symmetries with ultrastrong coupling and directional nonreciprocity at room temperature. These experimental results are reproduced well via numerical simulations and theoretical consideration. Ultrastrong coupling is due to a direct coupling between microwave photons confined in the chiral meta-atom and magnons in the magnetic meta-atom. Our results reveal a crucial step in identifying cavity-free, deepstrongly-coupled, nonreciprocal magnon-polaritons for hybrid quantum systems and quasi-particle “chemistry”using metamaterials.
@@
Introduction – Condensed matter physics is a field of various quasi-particles with elementary excitations. Integrating different quasi-particles has led to novel and intriguing properties of matters and opened an essential pathway toward frontiers in physics Kockum2019 ; Diaz2019 . A magnon, a quantized elementary excitation in magnets, can be transformed to magnon-polariton when it is coupled to alternating current (AC) electromagnetic fields, namely photons. The magnon-polariton has gained attention particularly in quantum information technology and spintronics Li2020 ; Harder2021 for applications to hybrid quantum systems Lachance-Quirion2019 .
@@
Magnons and photons are coherently coupled – strong coupling – if the coupling constant, $g$ is larger than the dissipation rate, $\kappa$, in a resonator and oscillator ($g>\kappa$), generating magnon-polariton. Here, $g$ is given by $g=(\omega_{+}-\omega_{-})/2$, where $\omega_{\pm}$ is the hybridized eigenmodes at the crossing point of the original modes. Such coherent coupling of magnons with photons with a frequency of $\omega/2\pi$ manifests itself at the anti-crossing of the dispersions. The anti-crossing is accompanied by Rabi-like level repulsion of $2g/2\pi$, which is proportional to a coupling ratio of $g/\omega$. While the strong coupling is characterized by $g>\kappa$, an ultrastrong magnon-photon coupling is represented by $g/\omega>0.1$ and $g>\kappa$, which is indispensable toward the quantum regime Bourcin2023 . Furthermore, the cavity-free magnon-polariton is a central topic in quantum optics as it allows control over individual quantum systems Bayer2017 ; Rajabali2022 . The cavity-free ultrastrongly-coupled magnon-polaritons have been demonstrated at microwave frequencies using magnetic heterostructures composed of superconductors Golovchanskiy2021PRAp ; Golovchanskiy2021SAd ; Ghirri2023 . They are, however, operated at cryogenic temperatures. Therefore, operating the ultrastrongly-coupled cavity-free magnon-polaritons at room temperature remains challenging.
@@
The ultrastrongly-coupled magnon-polariton is inherently nonreciprocal owing to the breaking of time-reversal symmetry by magnetization or magnetic fields Harder2018SSP . The nonreciprocity results in a transmission amplitude (and phase) difference depending on the propagation direction and polarization of photons. However, such a directionally nonreciprocal optical phenomenon has yet to be observed in ultrastrongly-coupled systems, except for dissipative Harder2018PRL or strong Zhang2020 magnon-photon coupling. Directionally nonreciprocal phenomena exist in natural chiral molecules Barron1984 ; Rikken1997 ; Vallet2001 , multiferroic materials Okamura2015 , and metamaterials Tomita2014 ; Tomita2017 ; Tomita2018 under the external direct current (DC) magnetic fields, as polarization-independent phenomena – magnetochiral (MCh) and optical magnetoelectric (ME) effects Tokura2018 ; Toyoda2019 . Although such nonreciprocal MCh metamaterials indicate the presence of room-temperature coherent couplings between magnons and photons at microwave frequencies Tomita2017 , the coupling mechanism remains unclear. The coupling mechanism should be therefore clarified by evaluating the coupling ratios of the magnon-polariton.
In this Letter, we demonstrate cavity-free ultrastrong-couplings of the directionally nonreciprocal magnon-polaritons in a single three-dimensional MCh metamolecule at room temperature. The coupling ration is evaluated to be $g/\omega>$ 0.2, which is achieved by direct interaction between microwave photons in the Cu chiral structure and magnons in the YIG cylinder. These experimental results are reproduced well via numerical simulations and theoretical consideration.
@@
Experimental setup – Figure 1(a) is an illustration of the MCh metamolecule. The MCh metamolecule consists of a polycrystalline yttrium-iron garnet (YIG) cylinder (as a magnetic meta-atom) inserted in a right-handed helix made of copper (Cu) (as a chiral meta-atom) placed in a microwave waveguide Tomita2017 . The Cu wire of 0.55 mm diameter is wounded four times with a 2.6 mm pitch to form the right-handed helix of a 2.55 mm outer diameter. The YIG cylinder is 2 and 15 mm in diameter and length, respectively.
@@
As in Fig. 1(b), the MCh metamolecule is set into a WR-90 waveguide that supports the TE10 mode with square flange adapters (Pasternack PE9804). The metamolecule is oriented along the $x$- or $z$-axis. See Fig. S1 in Supplemental Material SM for AC magnetic and electric fields of the TE10 mode in the WR-90 waveguide. An external DC magnetic field, $\mu_{0}H_{\rm ext}$, up to approximately 540 mT is applied in the +$z$ direction by an electromagnet. The waveguide is connected to a vector network analyzer (VNA) (Rohde & Schwarz ZVA67) with a microwave input power of 0 dBm (1mW). The $S$ parameters are measured using VNA. The $S_{21}$ indicates a complex transmission coefficient from port 1 to 2, while $S_{12}$ represents that from port 2 to 1. All measurements are carried out at room temperature.
@@
Experimental results – Figure 1(c) illustrates the transmission amplitude $|S_{21}|$ spectra of the $x$-axis-oriented MCh metamolecule from 7 to 14 GHz under varying $\mu_{0}H_{\rm ext}$ (0 - 543 mT). Without the DC magnetic field ($\mu_{0}H_{\rm ext}=0$), two sharp dips are observed: one at 9.34 GHz (A) and the other at 12.36 GHz (B). Based on numerical simulation using COMSOL Multiphysics of the metamolecule under $\mu_{0}H_{\rm ext}=0$, we confirm that a dip at the lowest frequency is caused by the third-order resonance, with two nodes in the Cu chiral meta-atom induced by the microwave photons with the AC magnetic field in the $x$-axis (See Fig. S2(a) and Movie S1 in Supplemental Material). The dip (A) at 9.34 GHz is thus traced back to the resonating photons, referred to as chiral cavity photons. Contrastingly, the dip (B) at the higher frequency is not caused by simple resonance but by a complex mode with tornado-like current distribution in the Cu chiral structure or Mie resonance by the displaced YIG cylinder (Figs. S2(a) and S2(c) and Movie S2 in Supplemental Material). Therefore, in this experiment, we focus on the third-order chiral cavity photon at approximately 9 GHz, which is labeled as (A).
@@
Applying the external DC magnetic field to the metamolecule in the +$z$-direction causes the chiral cavity photon mode (A) to shift slightly to a lower frequency at $\mu_{0}H_{\rm ext}$ = 61 mT and then turn to a higher frequency with increasing $\mu_{0}H_{\rm ext}$. The major blue shift of the chiral cavity photon mode is attributed to the magnetic permeability change of the YIG cylinder by the applied $\mu_{0}H_{\rm ext}$, which is the evidence of the interaction between the YIG magnetic meta-atom and the microwave photon in the Cu chiral meta-atom. When $\mu_{0}H_{\rm ext}$ = 312 mT, a slight dip appears at approximately 8 GHz labeled as (C) in Fig. 1(c). The slight dip keeps moving to a higher frequency with an increase in $\mu_{0}H_{\rm ext}$ up to 543 mT; therefore, the dip is caused by ferromagnetic resonance, i.e., a magnon, in the YIG magnetic meta-atom. When $\mu_{0}H_{\rm ext}$ = 339 mT, the magnon mode (C) approaches the chiral cavity photon mode (A). Simultaneously, another sharp dip appears at 7.19 GHz, denoted by (A’), and shifts to a higher frequency with a further increase in $\mu_{0}H_{\rm ext}$. This indicates the hybridization of chiral cavity photon modes (A) and (A’) with the magnon mode (C), demonstrating the coherent coupling between chiral cavity photons and magnons in the metamolecule.
@@
Figure 2: Two dimensional plot of measured and simulated magnon-polariton dispersions. Measured 2D plots of (a) transmission amplitude $|S_{21}|$ and (b) amplitude difference $|S_{21}|-|S_{12}|$ of $x$-axis-oriented metamolecule as functions of applied external DC magnetic field, $\mu_{0}H_{\rm ext}$, and frequency. (c) and (d) are 2D plots numerically simulated using COMSOL Multiphysics, corresponding to (a) and (b), respectively.
@@
In Fig. 2(a), $|S_{21}|$ is plotted two-dimensionally as functions of $\mu_{0}H_{\rm ext}$ of 0 - 543 mT (horizontal axis) and of frequency of 7 - 14 GHz (vertical axis) to evaluate magnon-photon coupling strength. The black color corresponds to a higher transmission in the $|S_{21}|$ spectra, while the yellow to white colors correspond to lower transmissions. Green dotted lines correspond to eye guides of the chiral cavity photon frequency and the magnon dispersion. The $|S_{21}|$ two-dimensional (2D) plot in Fig. 2(a) highlights, at $\mu_{0}H_{\rm ext}\sim$ 350 mT, the anti-crossing between the chiral cavity photon mode (A) with $\omega_{a}/2\pi$ = 9.34 GHz and magnon mode (C). The anti-crossing Rabi-like level splitting of chiral cavity photon modes (A) and (A’) is $g/\pi$ = 4.19 GHz at $\mu_{0}H_{\rm ext}$ = 367.7 mT, resulting in a coupling ratio of $g/\omega_{a}$ = 0.22. Figure 2(a) demonstrates that magnons in the YIG cylinder and photons in the Cu chiral structure in the MCh metamolecule are ultra-strongly coupled at room temperature. The metamolecule orientation can change the coupling ratio. Figures S3(a), S3(b), and S3(c) in Supplemental Material indicate that the $z$-axis-orientated metamolecule has a smaller coupling. The coupling ratio is thus relevant to the electromagnetic mode.
@@
Figures S4(a) and S4(b) in Supplemental Material present the directional transmission amplitude difference, $|S_{21}|-|S_{12}|$, and phase difference, $\arg S_{21}-\arg S_{12}$, spectra of the MCh metamolecule in the $x$-axis orientation, respectively. Using Fig. S4(a), a 2D plot of experimentally observed $|S_{21}|-|S_{12}|$ is drawn in Fig. 2(b). The red color corresponds to $|S_{21}|>|S_{12}|$, while the blue color corresponds to $|S_{21}|<|S_{12}|$. The chiral cavity photon mode (A) above the magnon mode line (C) demonstrates $|S_{21}|<|S_{12}|$, while the mode (A’) below the magnon mode (C) represents $|S_{21}|>|S_{12}|$. This is characteristic of the directional nonreciprocity of ultrastrongly-coupled magnon-polariton.
@@
Numerical simulation results – Simultaneous breaking of time-reversal and space-inversion symmetries by the coexistence of magnetism and chirality in the metamolecule is the primary origin of the directional nonreciprocity observed in Fig. 2(b). An alternative and artificial origin could be a slight displacement of the metamolecule from the waveguide center or a misalignment of the metamolecule from the $x$-axis because similar nonreciprocal signals have been observed by an asymmetrical microstrip line on a ferrite substrate Ueda2006 and a displaced magnetic material on a coplanar waveguide Kodama2018 . In this study, numerical simulation indicates that the dip (B) observed at a higher frequency is caused by the Mie resonance of the displaced YIG cylinder (Figs. S2(a) and S2(c) in Supplemental Material). Therefore, we conduct numerical simulations of the metamolecule under various magnetic fields using COMSOL Multiphysics to exclude such an artificial origin. The Supplemental Material details the numerical simulation method.
@@
Figure 2(c) presents a numerically simulated 2D plot of $|S_{21}|$ as functions of $\mu_{0}H_{\rm ext}$ of 200 - 600 mT (horizontal axis) and of 7 - 14 GHz frequency (vertical axis). Level repulsion due to the interaction between chiral cavity photons (D and D’) and magnons (E) is reproduced. The simulated coupling ratio of 0.15 derived from the level splitting of 3.21 GHz at $\mu_{0}H_{\rm ext}$ = 410 mT is smaller than that observed experimentally in Fig. 2(a) because the magnetization dynamics in the YIG cylinder is linearized in the numerical simulation. Nevertheless, the simulated 2D plot of $|S_{21}|-|S_{12}|$ in Fig. 2(d) has successfully reproduced the directional nonreciprocity of the magnon-polaritons. The metamolecule has negligible dissipation and is regarded as a Hermitian system. As shown in Fig. 2(d), removal of the misalignment and displacement of the metamolecule in the simulation verifies that the origin of directionally nonreciprocal signals is the magnon-polariton under the simultaneous breaking of time-reversal and space-inversion symmetries in the metamolecule.
@@
Theoretical considerations – The coupling ratio of the magnon-polariton is quantitatively reproduced by considering the magnetization dynamics. The experimental system is the forced oscillation of magnon-polaritons driven by the input microwaves; nonetheless, a simpler damped oscillation model is assumed to evaluate the coupling ratio of the magnon and chiral cavity photons. As in Fig. 3(a), $h_{\rm n,x}$ represents an AC magnetic field in the $x$-direction due to the Faraday electromagnetic induction in the chiral meta-atom – chiral cavity photons. Here, the integer $n$ denotes the mode-index of the chiral cavity photons. The AC magnetic field $h_{\rm n,x}$ excites magnons ($m_{\rm x}$) in the YIG cylinder. This leads to direct coupling between magnons and chiral cavity photons, generating ultrastrongly-coupled magnon-polaritons in the metamolecule. Recall that the numerical simulation using COMSOL indicates that the third-order resonance with $n$ = 3 ($h_{\rm 3,x}$) appears at approximately 10 GHz. Thus, the magnon-polaritons can be described as a coupled oscillator model that considers a uniform magnon mode and two chiral photon modes with $n$ = 3 and 5. The Supplemental Material presents the calculation details.
@@
Figure 3: Theoretical consideration including magnetization dynamics. (a) Schematic of MCh metamolecule composed of a Cu chiral structure involving a YIG cylinder. The metamolecule’s YIG cylinder is oriented to the $x$-axis. (b) and (c) Rabi-like oscillations of the chiral cavity photons ($h_{\rm n,x}$) and magnon ($m_{\rm x}$) under $\mu_{0}H_{\rm ext}$ = 0.39 T ($n$ = 3) and $\mu_{0}H_{\rm ext}$ = 0.50 T ($n$ = 5), respectively. (d) and (e) Corresponding Fourier spectra of chiral cavity photons ($h_{\rm n,x}$) for 9.34 GHz and 12.36 GHz, respectively.
@@
Figures 3(b) and 3(c) are calculated damped oscillations of the chiral cavity photons ($h_{\rm n,x}$) and the magnons ($m_{x}$) under $\mu_{0}H_{\rm ext}$ = 0.39 T for $n$ = 3 and under $\mu_{0}H_{\rm ext}$ = 0.50 T for $n$ = 5, respectively. We assume the higher frequency mode to be the fifth-order resonance. The time evolutions in Figs. 3(b) and 3(c) correspond to Rabi-like oscillations – the amplitude exchange between one chiral cavity photon (red or green) and a magnon (blue). A minimal complexity in the Rabi-like oscillations is attributed to interactions among three modes (one magnon and two chiral cavity photons) in the metamolecule.
@@
Figures 3(d) and 3(e) correspond to the Fourier spectra of the time evolutions, in which three peaks represent coupled modes among one magnon and two chiral cavity photons, i.e., magnon-polariton modes with the Rabi-like frequency splitting. According to these spectra, the evaluated couplings are $g_{3}/\pi$= 3.75 GHz and $g_{5}/\pi$ = 3.93 GHz, resulting in the coupling ratios of $g_{3}/\omega_{3}$ = 0.20 and $g_{5}/\omega_{5}$= 0.16, respectively. Here, the subscripts in $g_{n}$ and $\omega_{n}$ mean the corresponding values of the $n$-th order. Based on the experimental results, $\omega_{3}/2\pi$ = 9.34 and $\omega_{5}/2\pi$ = 12.36 GHz are assumed for this evaluation. The computed $g_{3}/\omega_{3}$ = 0.20 is consistent with the experimentally obtained $g/\omega_{\rm a}$ = 0.22, indicating that magnon-polariton in the MCh metamolecule is in the ultrastrong coupling regime. Moreover, $g_{3}/\omega_{3}$ = 0.20 and $g_{5}/\omega_{5}$ = 0.16 indicate that the coupling at a lower frequency is much stronger than that at a higher frequency.

### v3

@@ Abstract
We experimentally demonstrate magnon-polaritons with ultrastrong coupling and directional nonreciprocity in a metamolecule lacking time-reversal and space-inversion symmetries at room temperature. These experimental results are reproduced well via numerical simulations and theoretical consideration. Ultrastrong coupling is due to a direct interaction of magnons in the magnetic meta-atom with microwave photons confined in the chiral meta-atom as a resonator. Our results reveal a crucial step in identifying deepstrongly-coupled and optically-moving magnon-polaritons for hybrid quantum systems, synthetic gauge fields, and quasi-particle “chemistry” using metamaterials.
@@
Introduction – Condensed matter physics is a field of various quasi-particles with elementary excitations. Integrating different quasi-particles has led to novel and intriguing properties of matters and opened an essential pathway toward frontiers in physics Kockum2019 ; Diaz2019 . A magnon, a quantized elementary excitation in magnets, can be transformed to magnon-polariton (MP) when it is coupled to alternating current (AC) electromagnetic fields, namely photons. The MP has gained attention particularly in quantum information technology and spintronics Li2020 ; Harder2021 for applications to hybrid quantum systems Lachance-Quirion2019 .
@@
Magnons and photons are coherently coupled – strong coupling – if the coupling constant, $g$ is larger than the dissipation rate, $\kappa$, in a resonator and oscillator ($g>\kappa$), generating MP. Here, $g$ is given by $g=(\omega_{+}-\omega_{-})/2$, where $\omega_{\pm}$ is the hybridized eigenmodes at the crossing point of the original modes. Such coherent coupling of magnons with photons having a frequency of $\omega/2\pi$ gives rise to the dispersion anti-crossing due to the hybridization. The anti-crossing is evaluated using Rabi-like level repulsion of $2g/2\pi$, which is proportional to a coupling ratio of $g/\omega$. While the strong coupling is characterized by $g>\kappa$, an ultrastrong magnon-photon coupling is represented by $g>\kappa$ and $g/\omega>0.1$, which is indispensable toward the quantum regime Bourcin2023 . The ultrastrongly-coupled MPs have been demonstrated at microwave frequencies using magnetic heterostructures composed of superconductors Golovchanskiy2021PRAp ; Golovchanskiy2021SAd ; Ghirri2023 . They are, however, operated at cryogenic temperatures. Therefore, operating the ultrastrongly-coupled MPs at room temperature remains challenging.
@@
The ultrastrongly-coupled MP is inherently nonreciprocal owing to the breaking of time-reversal symmetry by magnetization or magnetic fields Harder2018SSP . With the additional breaking of the space-inversion symmetry, the nonreciprocity results in a transmission amplitude (and phase) difference depending on the propagation direction of photons. However, such a directionally-nonreciprocal optical phenomenon has yet to be observed in ultrastrongly-coupled systems, except for dissipative Harder2018PRL or strong Zhang2020 magnon-photon coupling. The directionally-nonreciprocal phenomena exist in natural chiral molecules Barron1984 ; Rikken1997 ; Vallet2001 , multiferroic materials Okamura2015 , and metamaterials Tomita2014 ; Tomita2017 ; Tomita2018 under the external direct current (DC) magnetic fields, as polarization-independent phenomena – magnetochiral (MCh) and optical magnetoelectric (ME) effects Tokura2018 ; Toyoda2019 . Although such nonreciprocal MCh metamaterials indicate the presence of room-temperature coherent couplings between magnons and photons at microwave frequencies Tomita2017 , the coupling mechanism remains unclear. The coupling mechanism should be therefore clarified by evaluating the coupling ratios of the MP.
In this Letter, we demonstrate experimentally and numerically ultrastrong couplings of the directionally-nonreciprocal MPs in a single three-dimensional MCh metamolecule at room temperature. The coupling ratio is evaluated to be $g/\omega>$ 0.2, which is achieved by direct interaction between magnons in the yttrium-iron garnet (YIG) cylinder and microwave photons in the Cu chiral structure as a resonator. The further enhancement of the coupling ratio can be achieved by a decrease in the frequency of the fundamental mode of the chiral resonance, bringing about the deepstrong-coupling for hybrid quantum systems. Moreover, the metamolecule including the chiral resonator paves a way to the directionally-nonreciprocal MPs in the free-space for synthetic gauge fields for photons. Furthermore, the synthesis of metamolecules and metamaterials by combining different meta-atoms leads to quasi-particle “chemistry”.
@@
Experimental setup – Figure 1(a) is an illustration of the MCh metamolecule. The MCh metamolecule consists of a polycrystalline YIG cylinder (as a magnetic meta-atom) inserted in a right-handed helix made of copper (Cu) (as a chiral meta-atom) Tomita2017 . The Cu wire of 0.55 mm diameter is wounded four times with a 2.6 mm pitch to form the right-handed helix of a 2.55 mm outer diameter. The YIG cylinder is 2 and 15 mm in diameter and length, respectively.
@@
As in Fig. 1(b), the MCh metamolecule is set into a WR-90 waveguide that supports the TE10 mode with square flange adapters (Pasternack PE9804). The metamolecule is oriented along the $x$-axis, which is parallel to the AC magnetic field in the WR-90 waveguide. See Fig. S1 in Supplemental Material SM for AC magnetic and electric fields of the TE10 mode in the waveguide. An external DC magnetic field, $\mu_{0}H_{\rm ext}$, up to approximately 540 mT is applied in the +$z$ direction by an electromagnet. The waveguide is connected to a vector network analyzer (VNA) (Rohde & Schwarz ZVA67) with a microwave input power of 0 dBm (1mW). The $S$ parameters are measured using VNA. The $S_{21}$ indicates a complex transmission coefficient from port 1 to 2, while $S_{12}$ represents that from port 2 to 1. All measurements are carried out at room temperature.
@@
Experimental results – Figure 1(c) illustrates the transmission amplitude $|S_{21}|$ spectra of the $x$-axis-oriented MCh metamolecule from 7 to 14 GHz under varying $\mu_{0}H_{\rm ext}$ (0 - 543 mT). Without the DC magnetic field ($\mu_{0}H_{\rm ext}=0$), two sharp dips are observed: one at 9.34 GHz (A) and the other at 12.36 GHz (B). Based on numerical simulation using COMSOL Multiphysics of the metamolecule under $\mu_{0}H_{\rm ext}=0$, we confirm that a dip at the lowest frequency is caused by the third-order resonance, with two nodes in the Cu chiral meta-atom induced by the microwave photons with the AC magnetic field in the $x$-axis (See Fig. S2(a) and Movie S1 in Supplemental Material SM ). The dip (A) at 9.34 GHz is thus traced back to the resonating photons, referred to as chiral resonant photons, because the chiral meta-atom functions as a resonator or sensitizer. Contrastingly, the dip (B) at the higher frequency is not caused by simple resonance but by a complex mode with tornado-like current distribution in the Cu chiral structure or Mie resonance by the displaced YIG cylinder (Figs. S2(a) and S2(c) and Movie S2 in Supplemental Material SM ). Therefore, in this experiment, we focus on the third-order chiral resonant photon at approximately 9 GHz, which is labeled as (A).
@@
Applying the external DC magnetic field to the metamolecule in the +$z$-direction causes the chiral resonant photon mode (A) to shift slightly to a lower frequency at $\mu_{0}H_{\rm ext}$ = 61 mT and then turn to a higher frequency with increasing $\mu_{0}H_{\rm ext}$. The major blue shift of the chiral resonant photon mode is attributed to the magnetic permeability change of the YIG cylinder by the applied $\mu_{0}H_{\rm ext}$, which is the evidence of the interaction between the YIG magnetic meta-atom and the microwave photon in the Cu chiral meta-atom. When $\mu_{0}H_{\rm ext}$ = 312 mT, a slight dip appears at approximately 8 GHz labeled as (C) in Fig. 1(c). The slight dip keeps moving to a higher frequency with an increase in $\mu_{0}H_{\rm ext}$ up to 543 mT; therefore, the dip is caused by ferromagnetic resonance, i.e., a magnon, in the YIG magnetic meta-atom. When $\mu_{0}H_{\rm ext}$ = 339 mT, the magnon mode (C) approaches the chiral resonant photon mode (A). Simultaneously, another sharp dip appears at 7.19 GHz, denoted by (A’), and shifts to a higher frequency with a further increase in $\mu_{0}H_{\rm ext}$. This indicates the hybridization of chiral resonant photon modes (A) and (A’) with the magnon mode (C), demonstrating the coherent coupling between chiral resonant photons and magnons in the metamolecule.
@@
Figure 2: Two dimensional plot of measured and simulated MP’s dispersions. Measured 2D plots of (a) transmission amplitude $|S_{21}|$ and (b) amplitude difference $|S_{21}|-|S_{12}|$ of $x$-axis-oriented metamolecule as functions of applied external DC magnetic field, $\mu_{0}H_{\rm ext}$, and frequency. (c) and (d) are 2D plots numerically simulated using COMSOL Multiphysics, corresponding to (a) and (b), respectively.
@@
In Fig. 2(a), $|S_{21}|$ is plotted two-dimensionally as functions of $\mu_{0}H_{\rm ext}$ of 0 - 543 mT (horizontal axis) and of frequency of 7 - 14 GHz (vertical axis) to evaluate magnon-photon coupling strength. The black color corresponds to a higher transmission in the $|S_{21}|$ spectra, while the yellow to white colors correspond to lower transmissions. Green dotted lines correspond to eye guides of the chiral resonant photon frequency and the magnon dispersion. The $|S_{21}|$ two-dimensional (2D) plot in Fig. 2(a) highlights, at $\mu_{0}H_{\rm ext}\sim$ 350 mT, the anti-crossing between the chiral resonant photon mode (A) with $\omega_{a}/2\pi$ = 9.34 GHz and magnon mode (C). The anti-crossing Rabi-like level splitting of chiral resonant photon modes (A) and (A’) is $g/\pi$ = 4.19 GHz at $\mu_{0}H_{\rm ext}$ = 367.7 mT, resulting in a coupling ratio of $g/\omega_{a}$ = 0.22. Figure 2(a) demonstrates that magnons in the YIG cylinder and photons in the Cu chiral structure in the MCh metamolecule are ultra-strongly coupled at room temperature. The coupling ratio can be varied by the metamolecule orientation. Figures S3(a), S3(b), and S3(c) in Supplemental Material SM indicate that the $z$-axis-orientated metamolecule has a smaller coupling. The coupling ratio is thus relevant to the electromagnetic mode.
@@
Figures S4(a) and S4(b) in Supplemental Material SM present the directional transmission amplitude difference, $|S_{21}|-|S_{12}|$, and phase difference, $\arg S_{21}-\arg S_{12}$, spectra of the MCh metamolecule in the $x$-axis orientation, respectively. Using Fig. S4(a), a 2D plot of experimentally observed $|S_{21}|-|S_{12}|$ is drawn in Fig. 2(b). The red color corresponds to $|S_{21}|>|S_{12}|$, while the blue color corresponds to $|S_{21}|<|S_{12}|$. The chiral resonant photon mode (A) above the magnon mode line (C) demonstrates $|S_{21}|<|S_{12}|$, while the mode (A’) below the magnon mode (C) represents $|S_{21}|>|S_{12}|$. This is characteristic of the directional nonreciprocity of ultrastrongly-coupled MP.
@@
Numerical simulation – Simultaneous breaking of time-reversal and space-inversion symmetries by the coexistence of magnetism and chirality in the metamolecule is the primary origin of the directional nonreciprocity observed in Fig. 2(b). An alternative and artificial origin could be a slight displacement of the metamolecule from the waveguide center or a misalignment of the metamolecule from the $x$-axis because similar nonreciprocal signals have been observed by an asymmetrical microstrip line on a ferrite substrate Ueda2006 and a displaced magnetic material on a coplanar waveguide Kodama2018 . Indeed in this study, numerical simulation indicates that the dip (B) observed at a higher frequency is caused by the Mie resonance of the displaced YIG cylinder (Figs. S2(a) and S2(c) in Supplemental Material SM ). Therefore, we conduct numerical simulations of the metamolecule under various magnetic fields using COMSOL Multiphysics to exclude such an artificial origin. The Supplemental Material SM details the numerical simulation method (see also reference Pozar therein).
@@
Figure 2(c) presents a numerically simulated 2D plot of $|S_{21}|$ as functions of $\mu_{0}H_{\rm ext}$ of 200 - 600 mT (horizontal axis) and of 7 - 14 GHz frequency (vertical axis). Level repulsion due to the interaction between chiral resonant photons (D and D’) and magnons (E) is reproduced. The simulated coupling ratio of 0.15 derived from the level splitting of 3.21 GHz at $\mu_{0}H_{\rm ext}$ = 410 mT is smaller than that observed experimentally in Fig. 2(a) because the magnetization dynamics in the YIG cylinder is linearized in the numerical simulation. Nevertheless, the simulated 2D plot of $|S_{21}|-|S_{12}|$ in Fig. 2(d) has successfully reproduced the directional nonreciprocity of the MPs. The metamolecule has negligible dissipation and is regarded as a Hermitian system. As shown in Fig. 2(d), removal of the misalignment and displacement of the metamolecule in the simulation verifies that the origin of directionally-nonreciprocal signals is the MP under the simultaneous breaking of time-reversal and space-inversion symmetries in the metamolecule.
@@
Theoretical consideration – The coupling ratio of the MP is quantitatively reproduced by considering the magnetization dynamics. The experimental system is the forced oscillation of MPs driven by the input microwaves; nonetheless, a simpler damped oscillation model is assumed to evaluate the coupling ratio of the magnon and chiral resonant photons in the following. As in Fig. 3(a), $h_{\rm n,x}$ represents an AC magnetic field in the $x$-direction due to the Ampère’s circuital law accompanied by the Faraday’s electromagnetic induction in the chiral meta-atom – chiral resonant photons. Here, the integer $n$ denotes the mode-index of the chiral resonant photons. The AC magnetic field $h_{\rm n,x}$ excites magnons ($m_{\rm x}$) in the YIG cylinder. This leads to direct coupling between magnons and chiral resonant photons, generating ultrastrongly-coupled MPs in the metamolecule. Recall that the numerical simulation using COMSOL indicates that the third-order resonance with $n$ = 3 ($h_{\rm 3,x}$) appears at approximately 10 GHz. Thus, the MPs can be described as a coupled oscillator model that considers a uniform magnon mode and two chiral resonant modes with $n$ = 3 and 5. The Supplemental Material SM presents the calculation details (see also reference Cao2015 therein).
@@
Figure 3: Theoretical consideration including magnetization dynamics. (a) Schematic of MCh metamolecule composed of a Cu chiral structure involving a YIG cylinder. The metamolecule’s YIG cylinder is oriented to the $x$-axis. (b) and (c) Rabi-like oscillations of the chiral resonant photons ($h_{\rm n,x}$) and magnon ($m_{\rm x}$) under $\mu_{0}H_{\rm ext}$ = 0.39 T ($n$ = 3) and $\mu_{0}H_{\rm ext}$ = 0.50 T ($n$ = 5), respectively. (d) and (e) Corresponding Fourier spectra of chiral resonant photons ($h_{\rm n,x}$) for 9.34 GHz and 12.36 GHz, respectively.
@@
Figures 3(b) and 3(c) are calculated damped oscillations of the chiral resonant photons ($h_{\rm n,x}$) and the magnons ($m_{x}$) under $\mu_{0}H_{\rm ext}$ = 0.39 T for $n$ = 3 and under $\mu_{0}H_{\rm ext}$ = 0.50 T for $n$ = 5, respectively. We assume the higher frequency mode to be the fifth-order resonance. The time evolutions in Figs. 3(b) and 3(c) correspond to Rabi-like oscillations – the amplitude exchange between one chiral resonant photon (red or green) and a magnon (blue). A minimal complexity in the Rabi-like oscillations is attributed to interactions among three modes (one magnon and two chiral resonant photons) in the metamolecule.
@@
Figures 3(d) and 3(e) correspond to the Fourier spectra of the time evolutions, in which three peaks represent coupled modes among one magnon and two chiral resonant photons, i.e., MP modes with the Rabi-like frequency splitting. According to these spectra, the evaluated couplings are $g_{3}/\pi$= 3.75 GHz and $g_{5}/\pi$ = 3.93 GHz, resulting in the coupling ratios of $g_{3}/\omega_{3}$ = 0.20 and $g_{5}/\omega_{5}$= 0.16, respectively. Here, the subscripts in $g_{n}$ and $\omega_{n}$ mean the corresponding values of the $n$-th order. Based on the experimental results, $\omega_{3}/2\pi$ = 9.34 and $\omega_{5}/2\pi$ = 12.36 GHz are assumed for this evaluation. The computed $g_{3}/\omega_{3}$ = 0.20 is consistent with the experimentally obtained $g/\omega_{\rm a}$ = 0.22, indicating that MP in the MCh metamolecule is in the ultrastrong coupling regime. Moreover, $g_{3}/\omega_{3}$ = 0.20 and $g_{5}/\omega_{5}$ = 0.16 indicate that the coupling at a lower frequency is much stronger than that at a higher frequency.